# Cuspidal edges on focal surfaces of regular surfaces with parabolic points

### v1

@@ 2.2.1. Limiting normal curvature and singular curvature
Let $f\colon U\to\mathbb{R}^{3}$ be a front and $\nu$ its unit normal vector. Assume that $f$ is a cuspidal edge at $p$. Then one can find a singular curve $\gamma(t)$ passing through $p=\gamma(0)$ and a null vector field $\eta$ around $p$. We set $\hat{\gamma}=f\circ\gamma$. Then along $\gamma$, we can define the following geometric invariants:
(2.6) $\kappa_{\nu}(t)$ $=\frac{\left\langle{\hat{\gamma}^{\prime\prime}(t)},{\nu(\gamma(t))}\right\rangle}{|\hat{\gamma}^{\prime}(t)|^{2}}=-\frac{\left\langle{\hat{\gamma}^{\prime}(t)},{\nu^{\prime}(\gamma(t))}\right\rangle}{|\hat{\gamma}^{\prime}(t)|^{2}},$
@@
where $\lambda$ is the signed area density function. We call $\kappa_{\nu}$ and $\kappa_{s}$ the limiting normal curvature and the singular curvature, respectively ([16, 22]).
It is known that $\kappa_{\nu}$ relates to boundedness of the Gaussian curvature near a cuspidal edge. In fact, Saji, Umehara and Yamada [22] showed that the Gaussian curvature of $f$ is bounded on a sufficiently small neighborhood of $p$ if and only if $\kappa_{\nu}\equiv 0$ along a singular curvature $\gamma$ passing through $p$ (see Figure 1).
(a) $\kappa_{\nu}=0$
(b) $\kappa_{\nu}\neq 0$
Figure 1. Cuspidal edges with vanishing $\kappa_{\nu}$ (left) and non-vanishing $\kappa_{\nu}$ (right).
On the other hand, $\kappa_{s}$ corresponds to convexity or concavity of a cuspidal edge (see Figure 2).
(a) $\kappa_{s}>0$
(b) $\kappa_{s}<0$
In particular, for the case where the Gaussian curvature $K$ is bounded near a cuspidal edge $p$, if $K$ takes non-negative values around $p$, then $\kappa_{s}$ takes non-positive values along the singular curve passing through $p$ ([22, Theorem 3.1]).
2.3. Rational order of functions
[…]
Let $h_{1},h_{2}\colon(\mathbb{R}^{2},p)\to\mathbb{R}$ be function germs. Let $Z=h_{2}^{-1}(0)$. Define a functions germ $g=h_{1}/h_{2}\colon(\mathbb{R}^{2}\setminus Z,p)\to\mathbb{R}$. Then the rational order $\operatorname{ord}_{p}{g}$ of $g$ at $p$ is defined by
$\operatorname{ord}_{p}{g}\coloneqq\operatorname{ord}_{p}{h_{1}}-\operatorname{ord}_{p}{h_{2}}.$
[…]
Example 2.7.
Let $h_{1},h_{2}\colon(\mathbb{R}^{2},0)\to\mathbb{R}$ be function germs given by
$h_{1}(u,v)=uv,\quad h_{2}(u,v)=u^{2}-v^{2}.$
[…]
$g(u,v)=\frac{h_{1}(u,v)}{h_{2}(u,v)}=\frac{uv}{u^{2}-v^{2}}.$
This cannot be extended as a bounded function on $(\mathbb{R}^{2},0)$. However, the rational order $\operatorname{ord}_{0}{g}$ of $g$ at 0 is
$\operatorname{ord}_{0}{g}=2-2=0.$
This implies that $g$ is rationally bounded at 0.
@@ 3. Parallel surfaces
by Rodrigues’ formula. Thus $\nu$ can be also a unit normal vector to $f^{t}$. Moreover, $f^{t}$ is a front for any $t\in\mathbb{R}\setminus\{0\}$ ([28]). By (2.1), the first and second fundamental forms of $f^{t}$ are
(3.3) $\operatorname{I}^{t}=(1-t\kappa_{1})^{2}Edu^{2}+(1-t\kappa_{2})^{2}Gdv^{2},\quad\operatorname{II}^{t}=(1-t\kappa_{1})\kappa_{1}Edu^{2}+(1-t\kappa_{2})\kappa_{2}Gdv^{2}.$
Therefore the Gaussian and mean curvatures for $f^{t}$ are
(3.4) $K^{t}=\frac{\kappa_{1}\kappa_{2}}{(1-t\kappa_{1})(1-t\kappa_{2})}=\frac{K}{1-2tH+t^{2}K},\quad H^{t}=\frac{H-tK}{1-2tH+t^{2}K},$
where $K$ and $H$ are the Gaussian and mean curvatures of $f$, respectively.
@@ 3.1. Singularities of parallel surfaces
Then $\tilde{\lambda}^{-1}(0)=S(f^{t})$. Thus $\tilde{\lambda}$ is an identifier of singularities of $f^{t}$. In addition, $\partial_{u}$ give a null vector field for $f^{t}$. By Fact 2.6, we have the following immediately.
Proposition 3.1 ([8, Theorems 3.4 and 3.5]).
@@
Corollary 3.2.
With the same situations in Proposition 3.1, we have the following.
  1. (1)
    If the parallel surface $f^{t}$ is one of a cuspidal edge, a swallowtail or a cuspidal butterfly at $p$, then $\operatorname{ord}_{p}\tilde{\lambda}=1$, where $\tilde{\lambda}$ is the function as in (3.5).
  2. (2)
    If $f^{t}$ is either a cuspidal lips or a cuspidal beaks at $p$, then $\operatorname{ord}_{p}\tilde{\lambda}(p)=2$.
Proof.
We take a curvature line coordinate system $(u,v)$. We assume that $p$ is the origin 0 of $\mathbb{R}^{2}$. Then we remark that
[…]
hold ([10]).
[…]
$\tilde{\lambda}(u,v)=uh_{1}(u,v)+vh_{2}(u,v).$
[…]
3.2. Behavior of the Gaussian curvature
We consider behavior of the Gaussian curvature $K^{t}$ of the parallel surface $f^{t}$ of a regular surface $f$ at a distance $t=1/\kappa_{1}(p)$. When $f^{t}$ has non-degenerate singularity, the following assertion holds.
Proposition 3.3.
Let $f\colon U\to\mathbb{R}^{3}$ be a regular surface and $f^{t}$ the parallel surface with $t=1/\kappa_{1}(p)$. Suppose that $f^{t}$ is one of a cuspidal edge, a swallowtail or a cuspidal butterfly at $p$. Then the Gaussian curvature $K^{t}$ is rationally bounded at $p$ if and only if $\kappa_{2}$ vanishes at $p$. Moreover, $K^{t}$ is bounded near $p$ if and only if $\kappa_{2}$ vanishes on $S(f^{t})$.
Proof.
By [16, Corollary C], $K^{t}$ is rationally bounded at $p$ if and only if the Gauss map $\nu$ of $f^{t}$ is singular at $p$. This is equivalent to the condition that $\kappa_{2}(p)=0$ because $\kappa_{1}(p)\neq 0$. Thus the first assertion holds. Moreover, $K^{t}$ is bounded near $p$ if and only if the second fundamental form of $f^{t}$ vanishes on $S(f^{t})$ ([22, Theorem 3.1]). By (3.3), it holds that
$\operatorname{II}^{t}=(1-t\kappa_{2})\kappa_{2}dv^{2}$
on $S(f^{t})$. Thus the second assertion holds. ∎
Proposition 3.4.
Let $f\colon U\to\mathbb{R}^{3}$ be a regular surface whose parallel surface $f^{t}$ is a cuspidal lips or cuspidal beaks at $p$, where $t=1/\kappa_{1}(p)$. Then the Gaussian curvature $K^{t}$ of $f^{t}$ is rationally bounded at $p$ if and only if $\operatorname{ord}_{p}\kappa_{2}=2$. Moreover, $K^{t}$ is bounded near $p$ if and only if the set of parabolic points $\kappa_{2}^{-1}(0)$ contains $S(f^{t})$.
Proof.
By the assumption and Corollary 3.2, we see that the function $\tilde{\lambda}$ as in (3.5) satisfies $\operatorname{ord}_{p}\tilde{\lambda}=2$. By (3.4) and $t=1/\kappa_{1}(p)$, the Gaussian curvature $K^{t}$ of $f^{t}$ can be written as
$K^{t}=\frac{\kappa_{1}\kappa_{2}}{t^{2}(\kappa_{1}(p)-\kappa_{1})(\kappa_{1}(p)-\kappa_{2})}=\kappa_{1}(p)^{2}\frac{\kappa_{1}\kappa_{2}}{\tilde{\lambda}(\kappa_{2}-\kappa_{1}(p))},$
it follows that the rational order $\operatorname{ord}_{p}K^{t}$ of $K^{t}$ satisfies $\operatorname{ord}_{p}K^{t}=0$ at $p$ if and only if $\operatorname{ord}_{p}\kappa_{2}=2$. This implies that $K^{t}$ is rationally bounded at $p$ if and only if $\operatorname{ord}_{p}\kappa_{2}=2$. The boundedness is also proved. ∎
@@ 3.3. The limiting normal curvature
We investigate the limiting normal curvature at a singular point of $f^{t}$.
Proposition 3.5.
Let $f\colon U\to\mathbb{R}^{3}$ be a regular surface without umbilic point and $\nu$ its unit normal vector. Suppose that the principal curvature $\kappa_{1}$ does not vanish on $U$. If the parallel surface $f^{t}$ with $t=1/\kappa_{1}(p)$ satisfies $\operatorname{rank}df^{t}_{p}=1$, then the limiting normal curvature $\kappa_{\nu}^{t}$ of $f^{t}$ satisfies
$\kappa_{\nu}^{t}=\frac{\kappa_{1}\kappa_{2}}{\kappa_{1}-\kappa_{2}}$
at $p$.
Proof.
We take a curvature line coordinates $(u,v)$ on $U$. We first consider the case where the parallel surface $f^{t}$ is a cuspidal edge at $p$. An identifier of singularities $\tilde{\lambda}$ for $f^{t}$ is given by (3.5). By Proposition 3.1, there exists a function $u(v)$ such that $\tilde{\lambda}(u(v),v)=0$. We set a curve $c(v)=(u(v),v)$, and define $\hat{c}(v)=f^{t}(u(v),v)$. Direct calculations show
@@
where $^{\prime}=d/dv$ and $\hat{\nu}(v)=\nu(u(v),v)$. Thus we have
@@
Hence we have the assertions. ∎
By this proposition, we get the following.
Corollary 3.6.
With the same assumptions in Proposition 3.5, the limiting normal curvature $\kappa_{\nu}^{t}$ of $f^{t}$ vanishes at $p$ if and only if $\nu$ is singular at $p$. Moreover, $\bm{e}_{2}$ is also singular at $p$.
Proof.
By Proposition 3.5, $\kappa_{\nu}^{t}(p)=0$ if and only if $\kappa_{2}(p)=0$ under the assumptions. On the other hand, $\nu$ is singular at $p$ if and only if $\kappa_{2}(p)=0$ in this case. Moreover, by (2.4), $\bm{e}_{2}$ is singular at $p$ if $\kappa_{2}(p)=0$. Thus we get the conclusions. ∎
@@ 4. Focal surfaces
Therefore $\bm{e}_{1}$ (resp. $\bm{e}_{2}$) as in (2.3) can be taken as a unit normal vector to $C_{1}$ (resp. $C_{2}$). Thus focal surfaces are frontals in this case. Moreover, the signed area density functions for $C_{1}$ and $C_{2}$ are
@@
(cf. [10, 20]). For a point $p\in S(C_{1})$ (resp. $p\in S(C_{2})$), we see that $\operatorname{rank}d(C_{1})_{p}=1$ (resp. $\operatorname{rank}d(C_{2})_{p}=1$). Thus one can take $\eta^{C_{1}}=\partial_{u}$ (resp. $\eta^{C_{2}}=\partial_{v}$) as a null vector field for $C_{1}$ (resp. $C_{2}$). Moreover, by Lemma 2.3, it holds that $(\bm{e}_{1})_{u}\neq 0$ (resp. $(\bm{e}_{2})_{v}\neq 0$) at $p\in S(C_{1})$ (resp. $p\in S(C_{2})$). Hence $C_{1}$ (resp. $C_{2}$) is a front at $p\in S(C_{1})$ (resp. $p\in S(C_{2})$) (see [28]). For singularities of focal surfaces, we have the following.
@@
We show the case of $C_{1}$. Let us take a curvature line coordinate system $(u,v)$. Assume that the parallel surface $f^{t}$ with $t=1/\kappa_{1}(p)$ of $f$ is a cuspidal edge at $p$. Then by Proposition 3.1, it holds that $(\kappa_{1})_{u}(p)\neq 0$. Thus $p\notin S(C_{1})$. This implies that the first assertion holds. Next, we assume that $f^{t}$ is one of a swallowtail, a cuspidal lips or a cuspidal beaks. Then by Proposition 3.1, it holds that
$(\kappa_{1})_{u}(p)=0,\quad(\kappa_{1})_{uu}(p)\neq 0,\quad(\kappa_{1})_{v}(p)\neq 0.$
On the other hand, since an identifier of singularities for $C_{1}$ can be taken as $\tilde{\lambda}^{C_{1}}=(\kappa_{1})_{u}$ and $\eta^{C_{1}}=\partial_{u}$ is a null vector field for $C_{1}$, we have $\eta^{C_{1}}\tilde{\lambda}^{C_{1}}=(\kappa_{1})_{uu}$, which does not vanish at $p$ by the assumption. Hence $C_{1}$ is a cuspidal edge at $p$ by Fact 2.6. Therefore we get the second assertion.
@@ 4.1. Geometric invariants at cuspidal edges of focal surfaces
We investigate geometric invariants of focal surfaces at cuspidal edges. Let $f\colon U\to\mathbb{R}^{3}$ be a regular surface parametrized by a curvature line coordinate system $(u,v)$. We then assume that $\kappa_{2}(p)=0$ and the focal surface $C_{1}$ is a cuspidal edge at $p\in U$. In this case, we shall consider geometric invariants called the limiting normal curvature $\kappa_{\nu}^{C_{1}}$ and the singular curvature $\kappa_{s}^{C_{1}}$ of $C_{1}$ at $p$.
By the assumption, $S(C_{1})=((\kappa_{1})_{u})^{-1}(0)$ and $(\kappa_{1})_{uu}(p)\neq 0$ hold. Thus there exists a regular curve $\gamma(v)=(u(v),v)$ ($|v-v_{0}|<\varepsilon$) with $\gamma(v_{0})=p$ such that $(\kappa_{1})_{u}(u(v),v)=0$ by the implicit function theorem. The curve $\gamma(v)$ is a singular curve for $C_{1}$. Thus the curve
@@
gives the singular locus of $C_{1}$ near $p$. We investigate the first and second differentials of $\hat{\gamma}$. By (4.2) and $(\kappa_{1})_{u}(u(v),v)=0$, it holds that
@@
$=\left(1-\frac{\kappa_{2}}{\kappa_{1}}\right)\Gamma^{1}_{22}f_{u}+\left(\frac{2\kappa_{2}(\kappa_{1})_{v}}{\kappa_{1}^{2}}-\frac{(\kappa_{2})_{v}}{\kappa_{1}}+\left(1-\frac{\kappa_{2}}{\kappa_{1}}\right)\Gamma^{2}_{22}\right)f_{v}$
$\hskip 142.26378pt+\left(\frac{2(\kappa_{1})_{v}^{2}}{\kappa_{1}^{3}}-\frac{(\kappa_{1})_{vv}}{\kappa_{1}^{2}}+\left(1-\frac{\kappa_{2}}{\kappa_{1}}\right)N\right)\nu$
@@
(4.8) $\hat{\gamma}^{\prime\prime}(v)=\left(1-\frac{\kappa_{2}(\gamma(v))}{\kappa_{1}(\gamma(v))}\right)\Gamma^{1}_{22}(\gamma(v))f_{u}(\gamma(v))+\left(\frac{2\kappa_{2}(\gamma(v))(\kappa_{1})_{v}(\gamma(v))}{\kappa_{1}(\gamma(v))^{2}}-\frac{(\kappa_{2})_{v}(\gamma(v))}{\kappa_{1}(\gamma(v))}+\left(1-\frac{\kappa_{2}(\gamma(v))}{\kappa_{1}(\gamma(v))}\right)\Gamma^{2}_{22}(\gamma(v))\right)f_{v}(\gamma(v))+\left(-\frac{\det\operatorname{Hess}(\kappa_{1})(\gamma(v))}{\kappa_{1}(\gamma(v))^{2}(\kappa_{1})_{uu}(\gamma(v))}+\frac{2(\kappa_{1})_{v}(\gamma(v))^{2}}{\kappa_{1}(\gamma(v))^{3}}+\left(1-\frac{\kappa_{2}(\gamma(v))}{\kappa_{1}(\gamma(v))}\right)N(\gamma(v))\right)\nu(\gamma(v)),$
where $\det\operatorname{Hess}(\kappa_{1})(\gamma(v))=(\kappa_{1})_{uu}(\gamma(v))(\kappa_{1})_{vv}(\gamma(v))-(\kappa_{1})_{uv}(\gamma(v))^{2}$.
For the limiting normal curvature, we get the following assertion.
Proposition 4.3.
Let $f\colon U\to\mathbb{R}^{3}$ be a regular surface parametrized by a curvature line coordinates $(u,v)$. Suppose that the focal surface $C_{1}$ of $f$ is a cuspidal edge at $p$ and $\kappa_{2}(p)=0$. Then the limiting normal curvature $\kappa_{\nu}^{C_{1}}$ at $p$ is calculated as
$\kappa_{\nu}^{C_{1}}(p)=-\frac{\kappa_{1}(p)^{3}(\kappa_{2})_{u}(p)G(p)}{\sqrt{E(p)}\left(\kappa_{1}(p)^{4}G(p)+(\kappa_{1})_{v}(p)^{2}\right)}$
Proof.

### v2

@@ 2.2.1. Limiting normal curvature and singular curvature
Let $f\colon U\to\mathbb{R}^{3}$ be a front and $\nu$ its unit normal vector field. Assume that $f$ is a cuspidal edge at $p$. Then one can find a singular curve $\gamma(t)$ passing through $p=\gamma(0)$ and a null vector field $\eta$ around $p$. We set $\hat{\gamma}=f\circ\gamma$. Then along $\gamma$, we can define the following geometric invariants:
(2.6) $\kappa_{\nu}(t)$ $=\frac{\left\langle{\hat{\gamma}^{\prime\prime}(t)},{\nu(\gamma(t))}\right\rangle}{|\hat{\gamma}^{\prime}(t)|^{2}}=-\frac{\left\langle{\hat{\gamma}^{\prime}(t)},{\nu^{\prime}(\gamma(t))}\right\rangle}{|\hat{\gamma}^{\prime}(t)|^{2}}\quad{\left({}^{\prime}=\frac{d}{dt}\right)},$
@@
where $\lambda$ is the signed area density function. We call $\kappa_{\nu}$ and $\kappa_{s}$ the limiting normal curvature and the singular curvature, respectively (see [13, 12, 20, 25]).
Let $f\colon U\to\mathbb{R}^{3}$ be a front. Let $p\in U$ be a rank one singular point of $f$. Assume that $p$ is a non-degenerate singular point other than a cuspidal edge. Then the limiting normal curvature $\kappa_{\nu}$ of $f$ at $p$ is given by
(2.7) $\kappa_{\nu}(p)=\lim_{t\to 0}\frac{\left\langle{\hat{\gamma}^{\prime\prime}(t)},{\nu(\gamma(t))}\right\rangle}{|\hat{\gamma}^{\prime}(t)|^{2}}=-\lim_{t\to 0}\frac{\left\langle{\hat{\gamma}^{\prime}(t)},{\hat{\nu}^{\prime}(t)}\right\rangle}{|\hat{\gamma}^{\prime}(t)|^{2}},$
where $\hat{\gamma}(t)=f(\gamma(t))$, $\hat{\nu}(t)=\nu(\gamma(t))$ and $\gamma(t)$ is the singular curve of $f$ through $p=\gamma(0)$ ([12, Proposition 2.9]). On the other hand, assume that $p$ is a degenerate singular point. Take a local coordinate system $(u,v)$ satisfying $f_{v}(p)=0$. Then the limiting normal curvature $\kappa_{\nu}(p)$ of $f$ at $p$ is defined by
(2.8) $\kappa_{\nu}(p)=\frac{\left\langle{f_{uu}(p)},{\nu(p)}\right\rangle}{|f_{u}(p)|^{2}}$
([12, (2.2)]). In this case, $\kappa_{\nu}(p)$ does not depend on the choice of local coordinates satisfying $f_{v}(p)=0$ ([12, Proposition 2.9]).
@@ 3. Parallel surfaces
by Weingarten formulas (2.3). Thus $\nu$ can be also a unit normal vector to $f^{t}$. Moreover, $f^{t}$ is a front for any $t\in\mathbb{R}\setminus\{0\}$ ([25]).
@@ 3.1. Singularities of parallel surfaces
Then $\tilde{\lambda}^{-1}(0)=S(f^{t})$. Thus $\tilde{\lambda}$ is an identifier of singularities of $f^{t}$. In addition, $\partial_{u}$ give a null vector field for $f^{t}$. By Fact 2.4, we have the following immediately.
Proposition 3.1 ([5, Theorems 3.4 and 3.5]).
@@
We remark that if $f^{t}$ ($t=1/\kappa_{1}(p)$) is a cuspidal lips, then $(\kappa_{1})_{uu}(p)\neq 0$. If we consider the parallel surface $f^{t}$ with $t=1/\kappa_{2}(p)$, $\tilde{\lambda}=\kappa_{2}-\kappa_{2}(p)$ is an identifier of singularities and $\partial_{v}$ is a null vector field for $f^{t}$. Thus one can obtain similar characterizations as Proposition 3.1.
@@ 3.2. The limiting normal curvature
We here consider the limiting normal curvature $\kappa_{\nu}^{t}$ of $f^{t}$ at a rank one singular point. In particular, we give the proof of Proposition 1.1.
Proof of Proposition 1.1.
We take a curvature line coordinates $(u,v)$. We show the case where $t=1/\kappa_{1}(p)$. We first consider the case where the parallel surface $f^{t}$ is a cuspidal edge at $p$. An identifier of singularities $\tilde{\lambda}$ for $f^{t}$ is given by (3.3). By Proposition 3.1, there exists a function $u(v)$ such that $\tilde{\lambda}(u(v),v)=0$. We set a curve $c(v)=(u(v),v)$, and define $\hat{c}(v)=f^{t}(u(v),v)$. Direct calculations show
@@
where $^{\prime}=d/dv$ and $\hat{\nu}(v)=\nu(u(v),v)$. By (2.6), we have
@@
Hence we have the assertions. For the case of $t=1/\kappa_{2}(p)$, one can show in a similar way. ∎
@@ 4. Focal surfaces
Therefore $\bm{e}_{1}$ (resp. $\bm{e}_{2}$) as in (2.2) can be taken as a unit normal vector to $C_{1}$ (resp. $C_{2}$). It is known that $C_{1}$ (resp. $C_{2}$) is a front ([25, Theorem 7.1.1]). Moreover, the signed area density functions for $C_{1}$ and $C_{2}$ are
@@
(cf. [8, 17]). For a point $p\in S(C_{1})$ (resp. $p\in S(C_{2})$), we see that $\operatorname{rank}d(C_{1})_{p}=1$ (resp. $\operatorname{rank}d(C_{2})_{p}=1$). Thus one can take $\eta^{C_{1}}=\partial_{u}$ (resp. $\eta^{C_{2}}=\partial_{v}$) as a null vector field for $C_{1}$ (resp. $C_{2}$). By Fact 2.4, we have the following.
Proposition 4.1 (cf. [25, Theorem 7.1.1]).
Let $f\colon U\to\mathbb{R}^{3}$ be a surface without umbilic point parametrized by a curvature line coordinate system $(u,v)$. Suppose that $\kappa_{1}$ (resp. $\kappa_{2}$) does not vanish on $U$. Then the focal surface $C_{1}$ (resp. $C_{2}$) is a cuspidl edge at $p\in U$ if and only if $(\kappa_{1})_{u}(p)=0$ and $(\kappa_{1})_{uu}(p)\neq 0$ (resp. $(\kappa_{2})_{v}(p)=0$ and $(\kappa_{2})_{vv}(p)\neq 0$).
By Propositions 3.1 and 4.2, the following assertions hold.
@@
We show the case of $C_{1}$. Let us take a curvature line coordinate system $(u,v)$. The first assertion holds by the previous discussion. Thus we show the second assertion. We assume that $f^{t}$ is one of a swallowtail, a cuspidal lips or a cuspidal beaks. Then by Proposition 3.1, it holds that $(\kappa_{1})_{u}(p)=0$ and $(\kappa_{1})_{uu}(p)\neq 0$. Thus by Proposition 4.1, the assertion follows.
@@ 4.1. Geometric invariants at cuspidal edges of focal surfaces
We investigate geometric invariants of focal surfaces at cuspidal edges. Let $f\colon U\to\mathbb{R}^{3}$ be a regular surface without umbilic points. Assume that $C_{i}$ ($i=1$ or $2$) is a cuspidal edge at $p\in U$. Then we consider the limiting normal curvature $\kappa_{\nu}^{C_{i}}$ and the singular curvature $\kappa_{s}^{C_{i}}$ of $C_{i}$ at $p$. In particular, we shall give the proof of Theorem 1.2.
Using a curvature line coordinate system $(u,v)$ for $f$, the limiting normal curvature and the singular curvature for $C_{i}$ obtained as follows.
Proposition 4.4.
Let $f\colon U\to\mathbb{R}^{3}$ be a regular surface without umbilic points. Assume that $f$ is parametrized by a curvature line coordinate system $(u,v)$.
  1. (1)
    If $C_{1}$ is a cuspidal edge at $p\in U$, then the limiting normal curvature $\kappa_{\nu}^{C_{1}}$ and the singular curvature $\kappa_{s}^{C_{1}}$ can be written as
    (4.4) $\kappa_{\nu}^{C_{1}}$ $=-\frac{\kappa_{1}^{3}(\kappa_{2})_{u}G}{\sqrt{E}((\kappa_{1}-\kappa_{2})^{2}\kappa_{1}^{2}G+(\kappa_{1})_{v}^{2})},$
    $\kappa_{s}^{C_{1}}$ $=-\operatorname{sgn}((\kappa_{1})_{uu}\kappa_{1}(\kappa_{1}-\kappa_{2}))\frac{\kappa_{1}^{3}\sqrt{G}X_{1}}{((\kappa_{1}-\kappa_{2})^{2}\kappa_{1}^{2}G+(\kappa_{1})_{v}^{2})^{3/2}}$
    at $p$, where $X_{1}$ is
    $X_{1}=(\kappa_{1}-\kappa_{2})\left((\kappa_{1}-\kappa_{2})\kappa_{1}\kappa_{2}G-\frac{\det\operatorname{Hess}(\kappa_{1})}{(\kappa_{1})_{uu}}\right)+(\kappa_{1})_{v}(2(\kappa_{1})_{v}+\Gamma^{2}_{22}(\kappa_{1}-\kappa_{2})-(\kappa_{2})_{v}).$
  2. (2)
    If $C_{2}$ is a cuspidal edge at $p\in U$, then the limiting normal curvature $\kappa_{\nu}^{C_{2}}$ and the singular curvature $\kappa_{s}^{C_{2}}$ can be written as
    (4.5) $\kappa_{\nu}^{C_{2}}$ $=-\frac{\kappa_{2}^{3}(\kappa_{1})_{v}E}{\sqrt{G}((\kappa_{1}-\kappa_{2})^{2}\kappa_{2}^{2}E+(\kappa_{2})_{u}^{2})},$
    $\kappa_{s}^{C_{2}}$ $=-\operatorname{sgn}((\kappa_{2})_{vv}\kappa_{2}(\kappa_{2}-\kappa_{1}))\frac{\kappa_{2}^{3}\sqrt{E}X_{2}}{((\kappa_{1}-\kappa_{2})^{2}\kappa_{2}^{2}E+(\kappa_{2})_{u}^{2})^{3/2}}$
    at $p$, where $X_{2}$ is
    $X_{2}=(\kappa_{2}-\kappa_{1})\left((\kappa_{2}-\kappa_{1})\kappa_{1}\kappa_{2}E-\frac{\det\operatorname{Hess}(\kappa_{2})}{(\kappa_{2})_{vv}}\right)+(\kappa_{2})_{u}(2(\kappa_{2})_{u}+\Gamma^{1}_{11}(\kappa_{2}-\kappa_{1})-(\kappa_{1})_{u}).$
Proof.
We show the case of $C_{1}$. By (4.3) and Proposition 4.1, $S(C_{1})=((\kappa_{1})_{u})^{-1}(0)$ and $(\kappa_{1})_{uu}(p)\neq 0$ hold. Thus there exists a regular curve $\gamma(v)=(u(v),v)$ ($|v-v_{0}|<\varepsilon$) with $\gamma(v_{0})=p$ such that $(\kappa_{1})_{u}(u(v),v)=0$ by the implicit function theorem. The curve $\gamma(v)$ is a singular curve for $C_{1}$, and the curve
@@
gives the singular locus of $C_{1}$ near $p$. By (4.2) and $(\kappa_{1})_{u}(u(v),v)=0$, it holds that
@@
$=\left(1-\frac{\kappa_{2}}{\kappa_{1}}\right)\Gamma^{1}_{22}f_{u}+\left(\frac{2\kappa_{2}(\kappa_{1})_{v}}{\kappa_{1}^{2}}-\frac{(\kappa_{2})_{v}}{\kappa_{1}}+\left(1-\frac{\kappa_{2}}{\kappa_{1}}\right)\Gamma^{2}_{22}\right)f_{v}+\left(\frac{2(\kappa_{1})_{v}^{2}}{\kappa_{1}^{3}}-\frac{(\kappa_{1})_{vv}}{\kappa_{1}^{2}}+\left(1-\frac{\kappa_{2}}{\kappa_{1}}\right)N\right)\nu$
@@
(4.10) $\hat{\gamma}^{\prime\prime}(v)$ $=\left(1-\frac{\kappa_{2}(\gamma(v))}{\kappa_{1}(\gamma(v))}\right)\Gamma^{1}_{22}(\gamma(v))f_{u}(\gamma(v))$
$\quad+\left(\frac{2\kappa_{2}(\gamma(v))(\kappa_{1})_{v}(\gamma(v))}{\kappa_{1}(\gamma(v))^{2}}-\frac{(\kappa_{2})_{v}(\gamma(v))}{\kappa_{1}(\gamma(v))}+\left(1-\frac{\kappa_{2}(\gamma(v))}{\kappa_{1}(\gamma(v))}\right)\Gamma^{2}_{22}(\gamma(v))\right)f_{v}(\gamma(v))$
$\quad\quad+\left(-\frac{\det\operatorname{Hess}(\kappa_{1})(\gamma(v))}{\kappa_{1}(\gamma(v))^{2}(\kappa_{1})_{uu}(\gamma(v))}+\frac{2(\kappa_{1})_{v}(\gamma(v))^{2}}{\kappa_{1}(\gamma(v))^{3}}+\left(1-\frac{\kappa_{2}(\gamma(v))}{\kappa_{1}(\gamma(v))}\right)N(\gamma(v))\right)\nu(\gamma(v)).$